OpenStudy (anonymous):

A light ray traveling from diamond into air has an angle of refraction of 56 degrees. A piece of glass is then placed right next to the diamond a) how will the angle of refraction change? b) Justify your answer by considering any changes to the speed of light. PLEASE HELP ME I DON'T UNDERSTAND. Also, could you give me a very detailed explanation, so i could understand, and explain me how you got your answer thank you. :)

OpenStudy (anonymous):

Hi, you just need snells law for this. Let me first explain with words then some maths! Firstly, refraction occurs as light speeds up or slows down as it moves across an interface between two materials of different density. The bigger the difference in density, the more it refracts. So, from air to diamond, you get a BIG refraction. If it was straight from glass into diamond, it would not refract as much. The light slows down, but not as much. Imagine if the two materials have the same density - the light wouldn't refract at all, and that explains why this is possible: http://www.youtube.com/watch?v=Wif_yGZcK18 Now, for the maths: \[n _{1}\sin \theta _{1} = n _{2}\sin \theta _{2}\] is snells law, where n is the refractive index (basically how dense to light) and theta is the angle of the light ray from the normal. the 1 and 2 are just 'material 1' and 'material 2'. http://www.math.ubc.ca/~cass/courses/m309-01a/chu/Fundamentals/snell.htm Hope this this helps, feel free to ask anything else